Joint profit maximizing output and Cournot model
Consider two firms facing the demand curve P=75-5C, where Q Q1 +Q2. The firms cost functions are Cl (Q1) = 15 +10Q1 and C2 (Q2)-10 + 20Q2. Suppose that both firms have entered the industry. What is the joint profit-maximizing level of output? How much will each firm produce? Combined, the firms will produce units of output, of which Firm 1 will produce units and Firm 2 will produceunits. (Enter a numeric response using a real number rounded to two decimal places.) What is each firms equilibrium output and profit if they behave noncooperatively? Use the Cournot model. If the firms compete, then Firm 1 will produce produce units of output. units of output and Firm 2 will
Answer #1

a) The demand curve of duopolist firms is

P = 75- 5Q= 75-5(Q1 +Q2 )..............(i)

The cost function of firm 1 is

C(Q1 ) = 15 +10Q1 ........(ii)

The cost function of firm 2 is

C(Q2 ) = 10+20Q2 .........(iii)

When both firms jointly produce the output their cost function will be

C(Q) = C(Q1 ) +C(Q2 ) = 15+10Q1 +10+20Q2 = 25+10Q+10Q2 ........(iv)

Therefore, marginal cost of both firms is

MC= d(C(Q)/dQ= 10

Also using eqtn (i), the marginal revenue of both firms is

MR = d(PQ)/dQ= 75-10Q

The profit maximizing condition of a firm is

                     MR =MC

        i.e., 75-10Q=10 or, Q=6.5

Therefore, profit maximizing output of both firms = 6.5 units and Firm 1 will produce 6.5/2=3.25 units and Firm 2 will produce 6.5/2=3.25 units

b) If both firms will compete with each other,

then the marginal profit of firm 1 (d\pi1 /dQ1) = MR1 -MC1 = 75-10Q1 -5Q2 - 10=65-10Q1 -5Q2 =0

and the marginal profit of firm 2 (d\pi2 /dQ2 ) = 75-10Q2 -5Q1 -20= 55-10Q2 -5Q1 =0

Therefore, cournot nash equilibrium conditions are,

         65-10Q1 -5Q2 =0

      or, 10Q1 +5Q2 =65 or, 2Q1 + Q2 =13............(v)

and 55-10Q2 -5Q1 =0

   or, 2Q2 +Q1 =11.......................(vi)

Solving (v) and (vi), we get Q1 = 5 and Q2 = 3

Therefore, if both firms compete, then firm 1 will produce 5 units and Firm 2 will produce 3 units.